I need help solving this problem: write the equation of the linear function that has a slope 2/3 and passes through the point (9,-1)
I think I use this equation: y-y1=m(x-x1) but not positive

Answer by stanbon(75887)

I need help solving this problem: write the equation of the linear function that has a slope 2/3 and passes through the point (9,-1)
I think I use this equation: y-y1=m(x-x1) but not positive
------------------
You can use that but it's faster to use y = mx + b
---
You know m = 2/3, and y = -1 when x = 9
Solve for "b":
------
-1 = (2/3)(9) + b
-1 = 6 + b
b = -7
-----
Equation:
y = (2/3)x -7
===================
Cheers,
Stan H.
===================